OpenStudy (qaisersatti):

pleease disregard the values i wrote the actual values are Vin 27 v R1 5 ohm r2 = 7 Ohms R3 12 Ohms A=6 and B=0.1

OpenStudy (qaisersatti):

1)Vo* (R3-(B*R1*R3)+R1+R2+A*R2)/(R3-B*R1*R3)=VIN Vo x(12-(0.1x5x12)+5+7+6x7)/(12-0.1x5x12)=27 Vo = 2.7V 2)i1=V0/(R3-B*R1*R3) = 2.7/(12-0.1x5x12)=0.45 3)vo = vth = 2.7V 4)Rth=Vth/Isc=2.7/0.45= 6 ohm

OpenStudy (qaisersatti):

please anyone confirm are these right answers

OpenStudy (qaisersatti):

anyone who can confirm me the values actually he has only one chance remaining to check....anyone please

OpenStudy (nurali):

1. 2.7 2. 0.45 3. 2.7 4. 10.8

OpenStudy (qaisersatti):

Nur ali u mean i1=2.7 vOUT=0.45 Vth=2.7 Rth=10.8 ???? is that what u mean brother..

OpenStudy (qaisersatti):

its vout= 02.7 I1 =0.45 vth 2.7 rth 10.8 thanks Nur

OpenStudy (nurali):

@andaluse 1. 5 2. 7.5 3. 7.5 4. 5

OpenStudy (nurali):

if correct please click best response